# Three dimensional end-to-end simulation for kilonova emission from a black-hole neutron-star merger

### v1

@@ 6 Summary and Discussions
We found more than a factor of 2 variation in the KN brightness depending on the viewing angle in our present BH-NS model. Such a variation in the brightness can induce the same degree of the systematic error in conducting the ejecta parameter estimation for the ejecta properties, such as the mass, velocity, and effective ejecta opacity, from the observational data. We should also note that we only focus on one single case of a BH-NS merger with the DD2 EOS, and it is not clear whether the property of KN is always the same for other setups of BH-NS mergers. For example, if the longitudinal opening angle of the ejecta is close to $2\pi$, the BH-NS KN can have the viewing-angle dependence in the brightness comparably strong to those of BNS mergers as we indeed see in the results of the axisymmetrized model (see Fig. 10). We further should note that uncomprehended systematic errors in the opacity and heating rate can induce large systematic errors in the ejecta parameter inference. In particular, the latter can be severe for KNe from BH-NS mergers since the uncertainty is more significant for the ejecta with low values of $Y_{e}$ ($<0.24$, see  Barnes et al. (2021); Zhu et al. (2021)). Hence, it is essential to consider that these systematic errors can significantly alter the results of the ejecta parameter inference, and those estimated values should be used with a great caution.
@@ Acknowledgements
Numerical computation was performed on Yukawa21 at Yukawa Institute for Theoretical Physics, Kyoto University and the Yamazaki, Sakura, Cobra, and Raven clusters at Max Planck Computing and Data Facility, and the Cray XC50 at CfCA of the National Astronomical Observatory of Japan. ND acknowledges support from Graduate Program on Physics for the Universe (GP-PU) at Tohoku University. This work was supported by Grant-in-Aid for Scientific Research of JSPS/MEXT (20H00158, 21H04997, 22KJ0317, 23H00127, 23H04894, 23H04900, 23H05432, and 23H01172) and JST FOREST Program (JPMJFR212Y).
@@ Appendix A Formulation
In this appendix, we describe the formulation of axisymmetric hydrodynamics equations in the spherical coordinates employed for the long-term evolution of ejecta. Throughout this appendix, the units of $c=1=G$ are employed where $G$ is the gravitational constant, unless otherwise mentioned.
@@ A.1 Basic equations
where $\rho$, $h$, $u^{\mu}$, and $P$ denote the rest-mass density, specific enthalpy, four velocity, and pressure, respectively. The equations of the Euler equation, energy equation, and continuity equation are given, respectively, by
@@
Equations (4), (5) and (6) are rewritten in the forms
$\partial_{t}S_{i}+\partial_{k}\left(S_{i}v^{k}+P\alpha\sqrt{\gamma}\delta^{k}_{i}\right)=$ $-S_{0}\partial_{i}\alpha+S_{k}\partial_{i}\beta^{k}$
$-\frac{1}{2}\alpha\sqrt{\gamma}S_{jk}\partial_{i}\gamma^{jk}-\frac{\alpha{\dot{\epsilon}_{\rm esc}}}{hw}S_{i},$ (7)
$\partial_{t}S_{0}$ $+\partial_{k}\left[S_{0}v^{k}+P\sqrt{\gamma}\left(v^{k}+\beta^{k}\right)\right]=$
$-\gamma^{ij}S_{i}\partial_{j}\alpha+\alpha\sqrt{\gamma}S_{ij}K^{ij}-\alpha\rho_{*}{\dot{\epsilon}_{\rm esc}},$ (8)
$\partial_{t}\rho_{*}+\partial_{k}\left(\rho_{*}v^{k}\right)=0,$ (9)
respectively. Here, $K_{ij}$ denotes the extrinsic curvature, and the other variables which newly appear in the above equations are defined as follows:
$\sqrt{\gamma}$ $={\rm det}\left(\gamma_{ij}\right),$
$\rho_{*}$ $=\rho w\sqrt{\gamma},$
$w$ $=\alpha u^{t},$
$S_{i}$ $=\rho_{*}{\hat{u}}_{i}=\rho_{*}hu_{i},$
$S_{0}$ $=\rho_{*}{\hat{e}}=\rho_{*}\left(hw-\frac{P}{\rho w}\right),$
$S_{ij}$ $=\rho hu_{i}u_{j}+P\gamma_{ij},$
$v^{i}$ $=\frac{u^{i}}{u^{t}}.$ (10)
@@
Here, the definitions for the new variables are given as follows:
@@ B.1 Time evolution of particles
Suppose we have a time series of three-velocity field $v^{i(n)}$, where $i$ is the index of the spatial coordinates, at a time slice $t=t^{(n)}$. Here, the velocity field is defined at the spatial points $(x^{1}_{j},x^{2}_{k},x^{3}_{l})$ discretely, where $j$, $k$, $l$ denote the grid points in the three-dimensional space. For a given spatial position of a particle $x^{i(n)}$ at the $n$th time slice, the coordinates of the particle in the $(n+1)$th time slice is solved with the so-called semi-implicit trapezoidal method (Fujibayashi et al., 2020a; Nishimura et al., 2015). In this method, we solve
@@
where the spatial coordinates of the particle are in the intervals $x^{1}_{j+1}$$x^{1}_{j}$, $x^{2}_{k+1}$$x^{2}_{k}$, and $x^{3}_{l+1}$$x^{3}_{l}$ at $t=t^{(n)}$. $\Delta t^{\prime}<\Delta t$ indicates that the particle position changes significantly within the time step $\Delta t$. In such a case, we perform several sub-steps between the $(n+1)$th and $n$th time slices. The velocity fields at a time $t$ between $t^{(n+1)}$ and $t^{(n)}$ are linearly interpolated from the two time slices.
@@ B.2 Initial position of the particle
where $\Delta\theta$ is the polar angle spacing of the particle-set locations and $\langle v^{r}\rangle$ is the average radial velocity of the particles located at a given time. In this way, the particles are located uniformly in space, i.e., in similar radial and lateral distances.
@@ B.3 Nucleosynthesis
In total, about 9600 tracer particles are generated with this method. Along the thermodynamical histories of the tracer particles, nucleosynthesis calculations are performed in the same manner as in Wanajo et al. (2022). The initial $Y_{e}$ for the nucleosynthesis calculation is taken from the final value of each tracer particle, which is that shown in Figs. 2, 4, and 5. We define the particles with $Y_{e}\leq 0.08$ as those of dynamical ejecta. They are found to have very similar nucleosynthesis results with only a small difference in actinide abundances. Therefore, for the particles of the dynamical ejecta, we assign the same elemental abundance and radioactive heating rate, which are the mass-weighted average of those in Wanajo et al. (2022). On the other hand, for the post-merger component ($Y_{e}>0.08$), we assign each particle with the abundance and heating rate obtained by the nucleosynthesis calculation along its thermodynamical history.

### v2

@@ 6 Summary and Discussions
We found more than a factor of 2 variation in the KN brightness depending on the viewing angle in our present BH-NS model. Such a variation in the brightness can induce the same degree of the systematic error in conducting the ejecta parameter estimation for the ejecta properties, such as the mass, velocity, and effective ejecta opacity, from the observational data. We should also note that we only focus on one single case of a BH-NS merger with the DD2 EOS, and it is not clear whether the property of KN is always the same for other setups of BH-NS mergers. For example, if the longitudinal opening angle of the ejecta is close to $2\pi$, the BH-NS KN can have the viewing-angle dependence in the brightness comparably strong to those of BNS mergers as we indeed see in the results of the axisymmetrized model (see Fig. 11). We further should note that uncomprehended systematic errors in the opacity and heating rate can induce large systematic errors in the ejecta parameter inference. In particular, the latter can be severe for KNe from BH-NS mergers since the uncertainty is more significant for the ejecta with low values of $Y_{e}$ ($<0.24$, see  Barnes et al. 2021; Zhu et al. 2021). Hence, it is essential to consider that these systematic errors can significantly alter the results of the ejecta parameter inference, and those estimated values should be used with a great caution.
@@ Acknowledgements
We thank the anonymous referee for the careful reading of our manuscript and valuable comments. Numerical computation was performed on Yukawa21 at Yukawa Institute for Theoretical Physics, Kyoto University and the Yamazaki, Sakura, Momiji, Cobra, and Raven clusters at Max Planck Computing and Data Facility, and the Cray XC50 at CfCA of the National Astronomical Observatory of Japan. ND acknowledges support from Graduate Program on Physics for the Universe (GP-PU) at Tohoku University. This work was supported by Grant-in-Aid for Scientific Research of JSPS/MEXT (20H00158, 21H04997, 22KJ0317, 23H00127, 23H04894, 23H04900, 23H05432, and 23H01172) and JST FOREST Program (JPMJFR212Y).
@@ Appendix A Formulation
In this appendix, we describe the formulation of hydrodynamics equations in the spherical coordinates employed for the long-term evolution of ejecta. Throughout this appendix, the units of $c=1=G$ are employed where $G$ is the gravitational constant, unless otherwise mentioned.
@@ A.1 Basic equations
where $\rho$, $h$, $u^{\mu}$, and $P$ denote the rest-mass density, specific enthalpy, four velocity, and pressure, respectively. The Euler equation, energy equation, and continuity equation are given, respectively, by
@@
Here, $K_{ij}$ denotes the extrinsic curvature, and the other variables which newly appear in the above equations are defined as follows:
$\sqrt{\gamma}$ $={\rm det}\left(\gamma_{ij}\right),$
$\rho_{*}$ $=\rho w\sqrt{\gamma},$
$w$ $=\alpha u^{t},$
$S_{i}$ $=\rho_{*}{\hat{u}}_{i}=\rho_{*}hu_{i},$
$S_{0}$ $=\rho_{*}{\hat{e}}=\rho_{*}\left(hw-\frac{P}{\rho w}\right),$
$S_{ij}$ $=\rho hu_{i}u_{j}+P\gamma_{ij},$
$v^{i}$ $=\frac{u^{i}}{u^{t}},$
@@ B.1 Time evolution of particles
Suppose we have a time series of three-velocity field $v^{i(n)}$, where $i(n)$ is the index of the spatial coordinates, at a time slice $t=t^{(n)}$. Here, the velocity field is defined at the spatial points $(x^{1}_{j},x^{2}_{k},x^{3}_{l})$ discretely, where $j$, $k$, $l$ denote the grid points in the three-dimensional space. For a given spatial position of a particle $x^{i(n)}$ at the $n$th time slice, the coordinates of the particle in the $(n+1)$th time slice is solved with the so-called semi-implicit trapezoidal method (Fujibayashi et al., 2020a; Nishimura et al., 2015). In this method, we solve
@@
where the spatial coordinates of the particle are in the intervals $x^{1}_{j+1}$$x^{1}_{j}$, $x^{2}_{k+1}$$x^{2}_{k}$, and $x^{3}_{l+1}$$x^{3}_{l}$ at $t=t^{(n)}$. $\Delta t^{\prime}<\Delta t$ indicates that the particle position changes significantly within the time step $\Delta t$. In such a case, we perform several sub-steps between the $(n+1)$th and $n$th time slices. The velocity fields at time $t$ between $t^{(n+1)}$ and $t^{(n)}$ are linearly interpolated from the two time slices.
@@ B.2 Initial position of the particle
where $\Delta\theta$ is the polar angle spacing of the particle-set locations and $\langle v^{r}\rangle$ is the average radial velocity of the particles located at given time. In this way, the particles are located uniformly in space, i.e., in similar radial and lateral distances.
@@ B.3 Nucleosynthesis
In total, about 9600 tracer particles are generated with this method. Along the thermodynamical histories of the tracer particles, nucleosynthesis calculations are performed in the same manner as in Wanajo et al. (2022). The initial value of $Y_{e}$ for the nucleosynthesis calculation is taken from the final value of each tracer particle, which is that shown in Figs. 3, 5, and 6. We define the particles with $Y_{e}\leq 0.08$ as those of dynamical ejecta. They are found to have very similar nucleosynthesis results with only a small difference in actinide abundances. Therefore, for the particles of the dynamical ejecta, we assign the same elemental abundance and radioactive heating rate, which are the mass-weighted average of those in Wanajo et al. (2022). On the other hand, for the other component ($Y_{e}>0.08$), we assign each particle with the abundance and heating rate obtained by the nucleosynthesis calculation along its thermodynamical history.